Name Computer Worked with: labpal data on 3. The vapor pressure of water at 25 c is o.0313 atm. Calculate the values of Kp and Kc at C the equilibrium: H20 (l) H20(g). 4. vinegar contains acetic acid, a weak acid that is partially dissociated in aqueous solution: CH3CooH(aq) H (aq) CH3Co2 (aq) a) Write the equilibrium constant expression for Kc. b) What is the value of Kc if the extent of dissociation of 1.0 M CH3cooH is 0.42%? Determination of an Equilibrium Constant
Answer #1

3)

H2O (l)   <—> H2O(g)

Kp = p(H2O(g))

   = 0.0313

T = 25.0 oC =(25.0 + 273) K= 298.0 K

Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant

Δ n = 1

Kp= Kc (RT)^Δn

0.0313 = Kc *(0.0821*298.0)^(1)

Kc = 1.279*10^-3

Answer:

Kp = 0.0313

Kc = 1.279*10^-3

I am allowed to answer only 1 question at a time